OpenStudy (anonymous):

If cosθ = 2/3 and 270° < θ 360° , then determine the exact value of (1/cotθ).

OpenStudy (apolloschariot):

Have you ever used a unit circle before?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

but only an intro

OpenStudy (bradely):

cosθ = 2/3 270° < θ <360° θ is in 4th quadrant tan θ is negative cosθ=side/hyp From the Pythagorean theorem opp = sqrt(3^2-2^2)=sqrt(5) 1/cotθ=tanθ=opp/side =sqrt5/2 source: http://www.mathskey.com/question2answer

OpenStudy (anonymous):

thank you :)